### v4

# Statistical Significance of Feature Importance Rankings

Abstract
Feature importance scores are ubiquitous tools for understanding the predictions of machine learning models. However, many popular attribution methods suffer from high instability due to random sampling. Leveraging novel ideas from hypothesis testing, we devise techniques that ensure the most important features are correct with high-probability guarantees. These are capable of assessing both the set of $K$ top-ranked features as well as the order of its elements. Given local or global importance scores, we demonstrate how to retrospectively verify the stability of the highest ranks. We then introduce two efficient sampling algorithms that identify the $K$ most important features, perhaps in order, with probability at least $1-\alpha$. The theoretical justification for these procedures is validated empirically on SHAP and LIME.
1 Introduction
Many machine learning (ML) algorithms have impressive predictive power but poor interpretability relative to simpler alternatives like decision trees and linear models. This tradeoff has motivated a wide body of work seeking to explain how black-box models make predictions [Belle and Papantonis, 2021]. Such work is essential for building trust in ML systems in areas like finance, healthcare, and criminal justice, in which the consequences of model misbehavior may be severe [Dubey and Chandani, 2022, Ferdous et al., 2020, Mandalapu et al., 2023, Berk, 2019]. Interpretable ML methods can also help develop understanding of complex processes, augmenting domain knowledge with new hypotheses.
To that end, feature importance scores quantify the contributions of input features to the model’s performance. These may explain model behavior at the resolution of an individual sample (local) or in aggregate (global). A wide range of methods have been proposed, detailed in Section 2.
The particular value of a feature’s importance score may be of less practical interest than its ranking: which features are the most important, and the order of these highlighted features. A scientist using a machine learning model to predict disease risk from a patient’s genetic profile will focus on the genes with the highest importance scores for further study, prioritizing the ranking of the metric over its specific values. Similarly, explanations often report only a small number of features in order not to overwhelm the user. LIME [Ribeiro et al., 2016], for example, explicitly regularizes to report a fixed number of features.
Figure 1: Instability of top-5 feature rankings, Adult Census Income dataset. StableSHAP, SPRT-SHAP, and Adj. S-LIME are our contributions, run at $\alpha=0.1$. The computational budget of StableSHAP and Shapley Sampling are the same.
Unfortunately, many standard attribution methods suffer from instability induced by random sampling. Rerunning the same procedure could yield different explanations for which features are the most important. Local methods such as SHAP and LIME require Monte Carlo sampling to calculate scores. Global scores like LOCO are calculated on some finite set of input samples, yet are used to infer across the entire population of samples. This inherent lack of reproducibility seriously undermines the credibility of these analyses [Yu and Kumbier, 2020].
To address this, we cast a vast number of local and global attributions into a unifying framework, based on properties of unbiasedness and asymptotic normality. For any method in this framework, we propose techniques to verify that the observed importance rankings are correct with high-probability guarantees. Amongst the highest-ranking features, these techniques assess the stability of both the top-$K$ set as well as their ordering relative to one another.
We provide retrospective tools that analyze the rankings of given feature attributions (Section 4). Then, we propose two sampling methods which ensure the $K$ highest-ranked features are correct with probability exceeding $1-\alpha$ (Section 5). The first of these methods leverages the Sequential Probability Ratio Test (SPRT) from Wald [1945] to sample continually. The second approach, more efficient for certain scores, iterates sample size calculations on ambiguously-ranked pairs. Applied to Shapley values, we refer to these algorithms as SPRT-SHAP and StableSHAP. Figure 1 highlights the improvement in stability of our sampling algorithms relative to baselines on SHAP and LIME.  [^1] Our code and experimental results are at https://github.com/jeremy-goldwasser/feature-rankings.
2 Feature Importance Scores
2.1 Local Feature Importance
A broad range of research has addressed the question of how best to attribute variable importance for individual predictions. SHAP [Lundberg and Lee, 2017] and LIME [Ribeiro et al., 2016] are amongst the most popular local attribution methods. Both are model-agnostic and entail random Monte Carlo computation. We provide a brief review of these two techniques, then introduce global feature importance.
2.1.1 SHAP
SHAP is a special case of the Shapley value, a seminal concept from game theory [Shapley, 1952]. Shapley values, and SHAP by extension, uniquely satisfy several reasonable desiderata for credit allocation. For any subset $S$ of the $d$ variables, define the value function $v:S\rightarrow\mathbb{R}$ and weighting kernel $w_{S}\leavevmode\nobreak\ =\leavevmode\nobreak\ \binom{d-1}{d-|S|-1}^{-1}$. The Shapley value for variable $j$ is
$\phi_{j}(v)=\frac{1}{d}\sum_{S\subseteq[d]\backslash\{j\}}w_{S}(v(S\cup\{j\})-v(S)).$ (1)
Within the context of local feature attribution, $v(S)$ is typically the prediction the model $\hat{f}$ would have made on input $x$ if it only had access to the features $x_{S}$ with indices in $S\subseteq[d]$. In theory, this may be obtained by refitting the model on those features [Strumbelj et al., 2009, Lipovetsky and Conklin, 2001] For computational convenience, however, it is common to sample the features in $S^{c}$, concatenate them onto $x_{S}$, and take the prediction $\hat{f}(x_{S},x_{S^{c}})$. The unknown features are usually sampled from their marginal distributions [Lundberg and Lee, 2017, Strumbelj and Kononenko, 2014], though conditional distributions $X_{S^{c}}|X_{S}$ may be used instead [Aas et al., 2021, Frye et al., 2020].
Computing the exact Shapley value (1) requires evaluating $v(S)$ for $O(2^{d})$ terms. When the dimensionality is large, this is computationally prohibitive, so approximation algorithms must be used instead. The basic approach is Shapley Sampling (Algorithm 7) [Strumbelj and Kononenko, 2014]. Rather than using all subsets of $[d]\backslash\{j\}$, this algorithm samples $n$ at random.
$\hat{\phi}_{j}(v)=\frac{1}{n}\sum_{i=1}^{n}v(S_{j}^{i}\cup\{j\})-v(S_{j}^{i}).$ (2)
This sample average is unbiased for the true Shapley value, meaning $\mathbb{E}[\hat{\phi}_{j}]=\phi_{j}$. Lundberg and Lee [2017] later proposed KernelSHAP, a more efficient estimator than Shapley Sampling. KernelSHAP obtains all $d$ Shapley values at once in a linear regression framework.
2.1.2 LIME
LIME explains an individual prediction by fitting an inherently interpretable model on samples generated around the input in question [Ribeiro et al., 2016]. These random samples are labeled with the prediction of the original model. The explanation model is often chosen to be a linear model, perhaps regularized for smoothness or sparsity. In that case, the regression coefficients provide natural feature importance scores.
The default explanation model in the lime package is a sparse linear model with $K$ nonzero coefficients. To select the features, one could choose the $K$ highest products of OLS coefficient and original feature value. Another option is to take the first $K$ features that enter the Lasso path with decreasing $\lambda$ [Tibshirani, 1996, Tibshirani and Taylor, 2011]. Finally, the authors propose K-Lasso, sequentially selecting features with Least Angle Regression (LARS) [Efron et al., 2004]. In this context, the ranking could be the order the features are selected, rather than their coefficient values.
2.2 Global Feature Importance
A number of methods summarize a model’s general behavior with global feature importance scores. The global scores $\hat{\phi}_{j}\;\forall j$ are calculated from a finite number of inputs $x_{i}$. They may be used as a proxy for $\phi_{j}$, the population score describing the model across the entire sample space $\mathcal{X}$.
Most global attribution methods are a sample average of scores on individual data points. For local score $s_{j}^{i}$ on feature $j$ of input $i$, these attributions are
$\hat{\phi}_{j}=\frac{1}{n}\sum_{i=1}^{n}s_{j}^{i}.$ (3)
One approach merely takes the average of local feature attribution scores. For example, SAGE is the mean LossSHAP score; this is a Shapley value based on the per-sample loss, rather than the prediction. A global Shapley method that does not require labels is the average absolute SHAP value [Lundberg and Lee, 2017, Lundberg et al., 2020]. Similarly, [Ribeiro et al., 2016] computes the mean absolute LIME score to present global explanations. Analogous strategies have been suggested for counterfactual explanations [Ley et al., 2022]. More broadly, van der Linden et al. [2019] proposes global importance scores that average the absolute value of any local metric.
Another subset of global scores taking the form (3) compares model performance before and after controlling for feature $j$. Performance is evaluated with some loss function $\mathcal{L}$ on labeled data points $\{(x_{i},y_{i})\}_{i=1}^{n}$. These methods modify either the data or model to produce some altered prediction $\tilde{y}_{ij}$. Their local scores take the form
$s_{j}^{i}=\mathcal{L}(y_{i},\hat{y_{i}})-\mathcal{L}(y_{i},\tilde{y}_{ij}).$
Here, we survey choices of $\tilde{y}_{ij}$ for a number of popular loss-based frameworks.
  • Permutation Importance [Breiman, 2001, Fisher et al., 2019]. Averaging across unseen samples, this approach takes predictions after permuting the values of feature $j$. Its original formulation, Mean Decrease in Accuracy (MDA), used out-of-bag training samples in random forests.
  • Conditional Variable Importance [Strobl et al., 2008]. Permutation Importance samples features from their marginal distribution, producing unlikely samples with potentially inaccurate predictions. Conditional Variable Importance instead samples each feature from its distribution conditional on the other features.
  • Leave-One-Covariate-Out [Lei et al., 2017, Verdinelli and Wasserman, 2023]. LOCO retrains the model without each feature $j$, then predicts $\tilde{y}_{ij}$ using all other features.
  • Permute-and-relearn. This technique retrains after imputing random values of a given feature from its marginal [Hooker et al., 2021] or conditional [Mentch and Hooker, 2016] distribution.
A separate strategy trains an interpretable surrogate model on the predictions of the black-box model [Molnar, 2022, Hinton et al., 2015]. When a linear model is used, its regression coefficients may be taken as importance scores in the same fashion as LIME.
3 Rank Verification Review
3.1 Stable Importance Rankings
A small body of work seeks to identify the most important features from estimated feature improtances. Neuhof and Benjamini [2024] established population rankings in retrospect via simultaneous confidence intervals, adjusted for multiple testing with procedures such as Holm’s method [Holm, 1979]. While valid at level $\alpha$, the use of such multiple testing corrections drastically reduces the power of this procedure. The p-values are inflated by a factor of $O(d^{2})$, the total number of pairwise comparisons.
A number of works propose sampling algorithms designed to stabilize the set of $K$ highest Shapley values [Narayanam and Narahari, 2008, Pliatsika et al., 2024, Chabrier et al., 2024, Kolpaczki et al., 2021, Kariyappa et al., 2024]. The latter two present probabilistic guarantees for their top-$K$ bandit algorithms, albeit in terms of unknown parameters based on the gaps between Shapley values. Kariyappa et al. [2024] only guarantees $\epsilon$-approximate solutions — allowing incorrect top-$K$ features to be included, so long as they are within $\epsilon$ of the true top $K$. The algorithms are further hindered by the use of loose bounds like the Bonferroni correction.
3.2 Gaussian Methodology
Given a set of random variables, it is often of interest to verify whether the highest observed value - or values - indeed matches the population ranking with high probability. The methods discussed in 3.1 do so imprecisely, relying on multiple testing adjustments, loose bounds, and unknown parameters. In contrast, more powerful verification methods have been studied under various probability distributions via selective inference Hung and Fithian [2019], Taylor and Tibshirani [2015].
Goldwasser et al. [2025] introduced the first selective inference-based methods to verify the ranks of Gaussian data with unequal variances. For $j\in[d]=\{1,\ldots,d\}$, we observe independent random variables $X_{j}\sim\mathcal{N}(\mu_{j},\sigma_{j}^{2})$, where $\sigma_{j}$ is known and $\mu_{j}$ is unknown. For notational convenience, sort data according to their order statistics, such that $X_{1}$ has the highest observed value, $X_{2}$ is second-largest, etc.
The primary objective is to verify the “winner" $X_{1}$ as the “best," meaning $\mu_{1}>\mu_{j}$ for all $j>1$. Its rank is verified upon rejecting the null hypothesis $H_{01}$ that $\mu_{1}$ is not the highest, conditioned on the selection event that $X_{1}$ wins.
$H_{01}:\bigcup_{j>1}\underbrace{\mu_{1}\leq\mu_{j}\ |\ X_{1}>\max_{k>1}X_{k}}_{H_{01j}}.$ (4)
To test this null, define $\bar{\mu}_{1j}=\frac{\sigma_{j}^{2}x_{1}+\sigma_{1}^{2}x_{j}}{\sigma_{1}^{2}+\sigma_{j}^{2}}$, $\bar{\sigma}_{1j}^{2}=\frac{\sigma_{1}^{4}}{\sigma_{1}^{2}+\sigma_{j}^{2}}$, and $\bar{\eta}_{1j}=\max(\bar{\mu}_{1j},\max_{k\neq 1,j}x_{k})$. This parameterizes a normal distribution with mean $\bar{\mu}_{1j}$ and variance $\bar{\sigma}_{1j}^{2}$, whose left tail has been truncated at $\bar{\eta}_{1j}$. For standard normal CDF $\Phi(\cdot)$, further define $p_{1j}$:
$p_{1j}=\frac{1-\Phi(\frac{x_{1}-\bar{\mu}_{1j}}{\bar{\sigma}_{1j}})}{1-\Phi(\frac{\bar{\eta}_{1j}-\bar{\mu}_{1j}}{\bar{\sigma}_{1j}})}.$ (5)
$p_{1j}$ is the tail mass of this truncated normal distribution above $x_{1}$. Goldwasser et al. [2025] shows it is a valid p-value for $H_{01j}$ (4). Furthermore, $H_{01}$ may be tested at level $\alpha$ with
$p_{1}^{*}=\max_{j>1}p_{1j}.$ (6)
The winner is verified when $p_{1}^{*}\leq\alpha$, or equivalently when all p-values $p_{1j}$ reject.
This result has useful extensions for top-$K$ rank verification, still without explicit multiple testing adjustments. These verify all ranks $i\leq K$ by rejecting nulls of the form
$H_{0i}:\bigcup_{j}\underbrace{\mu_{i}\leq\mu_{j}\ |\ X_{i}>\max_{k}X_{k}}_{H_{0ij}}.$ (7)
Procedure 1 (Top-$K$ Ranks). This outputs an integer $K\geq 0$ such that the probability that the top-$K$ rankings are correct is at least $1-\alpha$. It iterates the test in Equation (6) on successive ranks until a failure to reject (Alg. 3).
First, test $H_{01}$ with $p_{1}^{*}=\max_{j>1}p_{1j}$. If $p_{1}^{*}\leq\alpha$, then test $H_{02}=\bigcup_{j>2}H_{02j}$ with $p_{2}^{*}=\max_{j>1}p_{2j}$. Continuing for $K\geq 0$ rejections, stop at the first failure to reject, wherein $p_{(K+1)j}>\alpha$ for some $j>K+1$.
Procedure 2 (Top-$K$ Set). This test evaluates whether the set of $K$ largest observed elements are guaranteed to have the highest means with probability exceeding $1-\alpha$ (Alg. 4). Unlike in Procedure 1, here $K$ is fixed a priori, and the orderining within the top $K$ does not matter.
Procedure 2 tests $\widetilde{H}_{0i}=\bigcup_{j>K}H_{0ij}$, i.e. that $\mu_{i}$ is not actually in the top $K$ set, with $\tilde{p}_{i}^{*}=\max_{j>K}p_{ij}$. When all $K$ nulls $\widetilde{H}_{0i}$ reject, the top-$K$ set is verified. This is equivalent to having $p_{ij}\leq\alpha$ for all $i\leq K$ and $j>K$.
We refer the reader to the original manuscript for the proofs.
4 Retrospective Score Verification
Under mild assumptions, most of the scores discussed in Section 2 are normally distributed and unbiased. Their variance can be well-approximated with simple techniques.
To see this, consider first attributions of the form $\hat{\phi}_{j}=\frac{1}{n}\sum_{i=1}^{n}s_{j}^{i}$, such as Shapley Sampling and LOCO. When the samples used to compute $\hat{\phi}_{j}$ are selected at random, $\hat{\phi}_{j}$ is unbiased for the population score $\phi_{j}$. Moreover, with sufficiently large $n$, the distribution of $\hat{\phi}_{j}$ converges to a normal distribution centered around $\phi_{j}$ by the central limit theorem. The sample variance $\hat{\sigma}_{j}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(s_{j}^{i}-\bar{s}_{j})^{2}$ is unbiased for $\sigma_{j}^{2}$, converging at an $O(n^{-1/2})$ rate.
In addition, the scores of local (LIME) and global surrogate methods are linear regression coefficients. Assuming normal errors, Ordinary Least Squares coefficients are unbiased, and their variance can be estimated with standard formulas. Similarly, KernelSHAP expresses Shapley values as linear regression coefficients, solving with a weighted least squares. Covert and Lee [2020] proved it is asymptotically normal with negligible bias. They also introduced a variance estimator, studied and improved upon by Goldwasser and Hooker [2023]. Finally, LIME with K-Lasso selects features with scaled correlations that are asymptotically normal [Zhou et al., 2021].
Further assume the importance scores $\hat{\phi}_{j}$ are independent. This certainly holds in some cases. For example, Shapley Sampling (2) uses different subsets $S_{j}^{i}$ for each feature, so the resulting attributions are independent. However, mild correlation may exist in methods like KernelSHAP, where the same data is used to estimate multiple scores. Nevertheless, our empirical results indicate the ensuing procedures are always valid, and in fact somewhat conservative. Appendix D evaluates the merits of alternative approaches which use correlated testing.
Main result. Let $\hat{\phi}_{1},\ldots,\hat{\phi}_{d}$ be a set of such feature importance scores. Assume they are independent, normal, unbiased, and with known oracle variance ($\hat{\sigma}_{j}^{2}=\sigma_{j}^{2}$). Then the procedures from Goldwasser et al. [2025] may be applied to verify the observed importance rankings.
Let $(\hat{k})$ denote the $k$th largest score, perhaps sorting $\hat{\phi}$ by absolute value. Procedure 1 (Alg. 3) yields some non-negative integer $K$ such that with probability at least $1-\alpha$,
$\phi_{(\hat{1})}>\ldots>\phi_{(\hat{K})}>\max_{\ell>K}\phi_{(\hat{\ell})}.$
Procedure 2 (Alg. 4) tests the stability of the top-$K$ set, for any user-defined $K\in\{1,\ldots,d-1\}$. When its test rejects at level $\alpha$, the set of $K$ most important features is correct, again with the same high-probability guarantee.
The test (6) may also be applied simultaneously, indicating whether the $k$th-ranked feature is indeed higher than all lower ranks for all $k$ of interest. Doing so is akin to the lm function in R, which tests the statistical significance of all coefficients in a linear model. Not all significant results necessarily hold at level $\alpha$ due to multiple testing; nevertheless, this provides a concise summary of which ranks are likely stable.
5 Stabilized Top-$K$ Algorithms
This section introduces algorithms that guarantee the $K$ highest-ranking features, and perhaps their relative ordering, are correct with probability at least $1-\alpha$, where $K$ and $\alpha$ are predetermined by the user. This is motivated by numerous applications in which the $K$ most important features are analyzed, e.g. Narayanam and Narahari [2008], Goli and Mohammadi [2022], Ghorbani and Zou [2020]. Again, features may be sorted via their absolute values.
The algorithms sample until Procedure 1 or 2 verifies the top-$K$ features. To guarantee validity, Section 5.1 uses modified hypothesis tests based on the SPRT [Wald, 1945], whereas Section 5.2 performs sample size calculations to efficiently obtain each attribution.
5.1 SPRT Approach
A naive top-$K$ ranking strategy would draw samples until Procedure 1 from Section 3.2 rejects for the $K$ highest feature importances. An analogous approach for the top-$K$ set samples until Procedure 2 rejects. However, doing so would not necessarily control the error rate at level $\alpha$. This is because standard hypothesis tests like the p-values $p_{ij}$ from Goldwasser et al. [2025] are not valid under optional stopping, when data is accumulated until the moment it indicates a significant result. This process inflates the Type I error rate because it permits the data to be tested multiple times without adjustment.
To address this, we modify Procedures 1 and 2 so the tests they conduct are valid under optional stopping. The canonical choice in this setting is the Sequential Probability Ratio Test [Wald, 1945]. After any number of samples have been drawn, SPRT computes the likelihood ratio
$T_{n}=\frac{\max_{\theta\in H_{1}}\mathbb{P}(X\ |\ \theta)}{\max_{\theta\in H_{0}}\mathbb{P}(X\ |\ \theta)}.$ (8)
Set Type I and II error rates $\alpha$ and $\beta$, respectively. The test accepts $H_{1}$ when $T_{n}\geq\frac{1-\beta}{\alpha}$, accepts $H_{0}$ when $T_{n}\leq\frac{\beta}{1-\alpha}$, and continues sampling otherwise.
In the context of feature importance rankings, we construct the SPRT likelihood ratio $T_{n}$ with the following theorem. The proof, in Appendix C, involves maximum likelihood estimation in a selective inference framework.
Theorem 1.
Assume $X_{j}\sim\mathcal{N}(\mu_{j},\sigma_{j}^{2})$ independently, where $\sigma_{j}$ is known. Let $\phi$ and $\Phi$ be the standard normal PDF and CDF, and recall the definitions of $\bar{\mu}_{1j}$, $\bar{\sigma}_{1j}$, and $\bar{\eta}_{1j}$ from Section 3.2. For any Type I and II error thresholds $\alpha$ and $\beta$, a valid SPRT test statistic for $H_{01j}:\mu_{1}\leq\mu_{j}\ |\ X_{1}>\max_{k>1}X_{k}$ is
$T_{1j}^{n}=\left[\frac{\phi(0)}{1-\Phi(\frac{\bar{\eta}_{1j}-x_{i}}{\bar{\sigma}_{1j}})}\right]\left[\frac{\phi(\frac{x_{1}-\bar{\mu}_{1j}}{\bar{\sigma}_{1j}})}{1-\Phi(\frac{\bar{\eta}_{1j}-\bar{\mu}_{1j}}{\bar{\sigma}_{1j}})}\right]^{-1}.$
If $T_{1j}^{n}\geq\frac{1-\beta}{\alpha}$ for all $j$ such that $X_{i}>X_{j}$, then verify $\mu_{i}$ as larger than all $\mu_{j}$.
We employ Theorem (1) to learn the correct top-$K$ rankings (Alg. 1) and set (Alg. 5) with probability $1-\alpha$. In essence, our algorithms run computation until Procedures 1 and 2 verify the top $K$. However, the procedures are modified from Section 3.2: here, they test each null $H_{0ij}$ with the likelihood ratio $T_{ij}^{n}$, rather than p-values $p_{ij}$ (5). By SPRT, these tests can be conducted at any $n$ without regard for optional stopping.
Algorithm 1 Rank Stability via SPRT (SPRT-SHAP)
Desired rankings $K>0$, error rate $\alpha\in[0,1]$, total sample budget $n_{max}$, samples between tests $n_{btwn}$
$\hat{\phi}_{1},\ldots,\hat{\phi}_{d}$ whose top-$K$ rankings are correct with probability $\geq 1-\alpha$
$n\leftarrow 0$
while $n<n_{max}$ do
     Generate $n_{btwn}$ new samples
     $n\leftarrow n+n_{btwn}$
     $\hat{\phi}_{n}\leftarrow$ Feature importances fit on all $n$ samples
     $\hat{\Sigma}_{n}\leftarrow$ Variances of all feature importances
     $K^{\prime}\leftarrow$ Procedure 1 (Alg. 3) on $\hat{\phi}_{n}$ $\hat{\Sigma}_{n}$, rejecting
          tests $H_{0ij}$ if $T_{ij}^{n}>\frac{1-\beta}{\alpha}$ (Thm. 1)
     if $K^{\prime}>=K$ then
         return $\hat{\phi}$, “Verified"
     end if
end while
return $\hat{\phi}$, “Failed to verify"
5.2 Resampling Approach
While valid, SPRT is fairly conservative, as it must hold for any number of samples. An alternative approach tests with the original, more powerful p-values $p_{ij}$ (5). To account for optional stopping, it throws out all data used to compute $\hat{\phi}_{i}$ and $\hat{\phi}_{j}$ when null $H_{0ij}$ fails to reject. As a result, subsequent tests are independent of previous results. To guarantee validity, the earlier assumptions of independence, normality, unbiasedness, and oracle variance are still necessary.
This resampling strategy is only more efficient on certain importance scores. Shapley Sampling, for example, computes each attribution $\hat{\phi}_{j}$ in isolation with separate subsets $S_{j}$ (2). Therefore when $H_{0ij}$ fails to reject at level $\alpha$, only $\hat{\phi}_{i}$ and $\hat{\phi}_{j}$ must be recomputed; data for all $k\neq i,j$ may be kept. This is not the case for KernelSHAP, in which the same set of samples is used to compute all $d$ attributions $\hat{\phi}_{j}$. In that case, removing the entire dataset each time a pairwise test fails to reject is not viable.
Similarly, resampling may be more efficient than SPRT on global attributions computed in isolation (e.g. average Shapley Sampling, loss-based methods), but not jointly (e.g. average KernelSHAP, surrogate methods). Thus, all attributions that may benefit from this approach take the form
$\hat{\phi}_{j}=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}s_{j}^{i}.$
When some pairwise null $H_{0ij}$ fails to reject, the goal is to recompute $\hat{\phi}_{i}$ and $\hat{\phi}_{j}$ with as few samples as possible for $H_{0ij}$ to subsequently reject. Here we present two methods to approximate this. Let $\tilde{\sigma}_{j}^{2}$ be the sample variance of $s_{j}^{i}$, so $\sigma_{j}^{2}=\text{Var}(\hat{\phi}_{j})=\frac{\tilde{\sigma}_{j}^{2}}{n_{j}}$. Goldwasser et al. [2025] show that when the runner-up has highest p-value, Equation (5) reduces to a Z-test at level $\alpha/2$. Assuming this occurs,
$\frac{\hat{\phi}_{i}-\hat{\phi}_{j}}{\sqrt{\frac{\tilde{\sigma}_{i}^{2}}{n_{i}}+\frac{\tilde{\sigma}_{j}^{2}}{n_{j}}}}<Z_{1-\alpha/2}.$ (9)
Suppose we want the new number of samples $n_{i},n_{j}$ to be the same $n^{\prime}$ for both features. Then solving Equation (9) for $n^{\prime}$ leads to the sample size
$n^{\prime}=\Big{(}\frac{Z_{1-\alpha/2}}{\hat{\phi}_{i}-\hat{\phi}_{j}}\Big{)}^{2}(\tilde{\sigma}^{2}_{i}+\tilde{\sigma}^{2}_{j}).$ (10)
Alternatively, we may want the sample size to scale with the variance. Intuitively, more samples should be used to stabilize highly variable features. Defining $n_{j}^{\prime}:=\frac{\tilde{\sigma}^{2}_{j}}{\tilde{\sigma}^{2}_{j}}n_{i}^{\prime}$ and solving for $n_{i}^{\prime}$ yields
$n_{i}^{\prime}=2\Big{(}\frac{Z_{1-\alpha/2}}{\hat{\phi}_{i}-\hat{\phi}_{j}}\Big{)}^{2}\tilde{\sigma}^{2}_{i},\ \ n_{j}^{\prime}=2\Big{(}\frac{Z_{1-\alpha/2}}{\hat{\phi}_{i}-\hat{\phi}_{j}}\Big{)}^{2}\tilde{\sigma}^{2}_{j}.$ (11)
These lower bounds estimate the minimum number of samples needed to obtain an anticipated significant result. To avoid narrowly missing the mark, it is reasonable to choose values of $n^{\prime}$ that exceed them by a small buffer, e.g. 10%. It is entirely possible that more optimal choices of $n_{i}^{\prime}$ and $n_{j}^{\prime}$ exist; we leave this as an open problem.
Algorithm 2 Rank Stability via Resampling (StableSHAP)
Desired rankings $K>0$, error rate $\alpha\in[0,1]$, per-feature sample budget $n_{max}$, intial per-feature samples $n_{init}$, buffer $c\geq 1$
$\hat{\phi}_{1},\ldots,\hat{\phi}_{d}$ whose top-$K$ rankings are correct with probability $\geq 1-\alpha$
ϕ^^italic-$\hat{\phi}\leftarrow$ Feature importances fit on $n_{init}$ samples
$\hat{\Sigma}\leftarrow$ Variances of all feature importances
while $n_{j}<n_{max}\;\forall j$ do
     $K^{\prime}\leftarrow$ Procedure 1 (Alg. 3) on $\hat{\phi}_{n}$ $\hat{\Sigma}_{n}$, rejecting
          tests $H_{0ij}$ if $T_{ij}^{n}>\frac{1-\beta}{\alpha}$ (Thm. 1)
     if $K^{\prime}>=K$ then
         return $\hat{\phi}$, “Verified"
     else
         $n_{i}^{\prime},n_{j}^{\prime}\leftarrow$ Est. samples to reject, Eq. (10) or (11).
         $n_{i}^{\prime},n_{j}^{\prime}\leftarrow\min(\lceil cn_{i}^{\prime}\rceil,n_{max}),\min(\lceil cn_{j}^{\prime}\rceil,n_{max})$
         $\hat{\phi}_{i},\hat{\phi}_{j}\leftarrow$ Attributions fit on $n_{i}^{\prime},n_{j}^{\prime}$ samples
         $\hat{\Sigma}_{i},\hat{\Sigma}_{j}\leftarrow$ Variances of feature importances
     end if
end while
return $\hat{\phi}$, “Failed to verify"
Running for the maximal number of samples is not guaranteed to yield $K$ rejections at the desired error tolerance. Nevertheless, it may suffice even when the anticipated requisite budget is higher, since Equations (10) and (11) rely on plug-in variance estimates.
6 Experiments
To test the efficacy of our retrospective and top-$K$ methods, we run them on standard estimates of SHAP values. These attributions describe neural network classifiers fit on the Adult Census Income, Portuguese Bank, BRCA, Wisconsin Breast Cancer, and German Credit datasets.
Further experiments demonstrate the efficiency of the resampling approach, as well as rank stability on LIME (Appendix E, F). Appendix G contains more detailed information on these experiments.
6.1 Retrospective Stability
We evaluated the retrospective tools in Section 4 across a range of datasets, attribution methods, and significance levels. To do so, we randomly selected 30 input samples from the test set of each benchmark dataset. On each sample, we ran Shapley Sampling and KernelSHAP 50 times. For each iteration, we identified the number of stable ranks (Procedure 1) and assessed the stability of the top-$5$ set (Procedure 2).
We then computed the family-wise error rate (FWER) of the ranking and set procedures on each input sample. The FWER is the fraction of iterations with an error in the ranking or set of supposedly stable features. Iterations that do not verify any rankings or the top-$K$ set are counted as error-free. We used the most common top-$K$ ranking as the ground truth; this is almost certainly correct, since the Shapley estimators are unbiased.
Table 1: Maximum error rate of retrospective rank (R) and set (S) procedures, across 30 samples. $K=5$ for set.
Shapley Sampling KernelSHAP
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$ $\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Dataset N D R S R S R S R S R S R S
Adult 32,561 12 2% 4% 6% 4% 10% 10% 4% 4% 4% 6% 8% 12%
Bank 45,211 16 2% 0% 4% 2% 12% 8% 2% 0% 4% 0% 8% 2%
BRCA 572 20 2% 2% 8% 4% 10% 8% 2% 2% 2% 4% 10% 6%
Credit 1,000 20 2% 0% 8% 4% 10% 6% 2% 0% 2% 0% 6% 2%
WBC 569 30 2% 2% 8% 6% 10% 8% 6% 2% 6% 4% 10% 10%
Table 1 reports the maximal error rates across the 30 samples. In all 60 settings, the 30 FWERs are at most $\alpha$, indicating these procedures successfully control the FWER. This supports the application of Goldwasser et al. [2025] in the context of feature importances.
Figure 2 shows the number of verified ranks on the German Credit dataset. The median rises from one to two as the tolerable error threshold $\alpha$ goes to 0.2. In the most extreme case, the ranking of the top six features is stable. However, a substantial fraction of inputs had only zero or one stable ranking, emphasizing the fragile reliability of Shapley rankings.
Figure 2: Retrospective Verifications of KernelSHAP.
6.2 Top-K Rank Verification
Empirical results demonstrate that our sampling algorithms identify the $K$ most important features with probability $1-\alpha$. Applied to Shapley estimation, we refer to these algorithms as SPRT-SHAP (1) and StableSHAP (2). SPRT-SHAP obtains Shapley values with KernelSHAP, whereas StableSHAP is based on Shapley Sampling.
Table 2 summarizes their performance across a range of datasets, desired ranks $K$, and significance levels $\alpha$. As in Section 6.1, it presents the highest FWER over 30 inputs, calculated over 50 runs. In all settings for which the algorithms converged within the given sample budget, the FWER was below $\alpha$.
Table 2: Maximum error rate of top-$K$ rank (R) and set (S) procedures, across 30 samples. NAs indicate the procedure does not reject all $K$ tests on the provided input samples.
StableSHAP SPRT-SHAP
$K=2$ $K=5$ $K=2$ $K=5$
$\alpha=0.1$ $\alpha=0.2$ $\alpha=0.1$ $\alpha=0.2$ $\alpha=0.1$ $\alpha=0.2$ $\alpha=0.1$ $\alpha=0.2$
Dataset R S R S R S R S R S R S R S R S
Adult 8% 2% 16% 16% 6% 6% 14% 14% 0% 0% 2% 0% NA 0% NA 8%
Bank 6% 2% 14% 0% 10% 10% 20% 16% 0% 0% 2% 2% NA 0% NA 2%
BRCA 6% 6% 14% 10% 10% 10% 20% 20% NA 0% NA 0% NA 0% NA 2%
Credit 4% 2% 8% 2% 4% 4% 12% 16% 0% 0% 0% 0% NA 0% NA 2%
WBC 0% 4% 10% 12% 6% 4% 20% 4% 0% 0% 4% 0% NA 0% 0% 0%
StableSHAP is highly adaptive to the significance level. Its empirical FWERs were closest to $\alpha$, getting up to a 10% FWER with $\alpha=0.1$ and 20% with $\alpha=0.2$. In contrast, SPRT-SHAP produced more stable top-$K$ rankings, if it converged at all. This matches our intuition that SPRT is more conservative, requiring a higher evidence threshold to reject.
Figure 3: StableSHAP Sample Allocation.
Moreover, StableSHAP allocates computation in a highly efficient manner. It adaptively focuses on the features whose rankings are both ambiguous and relatively high. Figure 3 displays its sample allocations for a given input from the Adult dataset, running 50 times at $\alpha=0.1$. StableSHAP uses only the initial 100 samples for the highest-ranked feature, Marital Status, as it wins by a wide margin. It also avoids precise estimation of ranks beyond the top $K$ and its runner-up, Capital Gain. The main exception, Occupation, has considerably higher variance, thereby raising its p-values (Figure 6).
In contrast, most existing Shapley algorithms allot the same budget for all features. Figure 5 in Appendix E compares the performance of StableSHAP and Shapley Sampling, given the same inputs and computational budget. By and large, StableSHAP demonstrates improved stability, with fewer misranked features.
Figure 4 visualizes the average runtime of StableSHAP and SPRT-SHAP. The methods share the same minimal and maximal budget, and explain the same data. While both often achieve top-$2$ rank stability with the initial number of samples, StableSHAP is generally more efficient. SPRT-SHAP often takes a larger number of samples to converge, as evidenced by its higher third quartiles. Overall, these sample budgets are similar to the defaults in the shap package (Appendix G.1).
Figure 4: StableSHAP vs SPRT-SHAP Efficiency.
7 Discussion
In this paper we present methods to obtain stable orderings of feature importance scores. For a user-defined error rate $\alpha$, our retrospective procedures verify the highest observed rankings and top-$K$ set. Our top-$K$ algorithms efficiently run computation until the population rankings are attained with high probability. Our statistical guarantees, contingent on normal assumptions or asymptotic arguments, are unanimously justified by empirical results.
Our methods can be used to rank Shapley values in any context, not specifically for feature attributions. Other use cases include feature selection [Cohen et al., 2007], federated learning [Liu et al., 2022], data valuation [Ghorbani and Zou, 2019], multi-agent RL [Li et al., 2021], and ensembling [Rozemberczki and Sarkar, 2021]. Outside of ML, they have been applied in fields as diverse as ecology [Haake et al., 2007], online advertising [Zhao et al., 2018], supply chain management [Xu et al., 2018], and financial portfolio optimization [Shalit, 2020].
In addition to their analytical utility, our top-$K$ methods run relatively efficiently. The resampling approach adaptively budgets computation towards the more important features. And SPRT for Shapley importances enables use of KernelSHAP, a highly efficient algorithm. While SPRT is fairly conservative due to its anytime-valid requirement, future work could use more relaxed procedures that restrict the number of potential rejection times. Testing procedures that do so have been proposed in the context of clinical trials, e.g. Pocock [1977], O’Brien and Fleming [1979], Peto et al. [1976].
Our methods provide concrete statistical guarantees on their rankings and selections, with higher power and more reasonable assumptions than prior work [Kariyappa et al., 2024, Kolpaczki et al., 2021, Zhou et al., 2021]. Moreover, they may be applied in conjunction with methods that stabilize the attributions themselves [Goldwasser and Hooker, 2023, Mitchell et al., 2022].
Appendix A Procedures 1 and 2
Algorithm 3 Procedure 1 [Goldwasser et al., 2025]
Significance level $\alpha$, ordered data $X_{1}>X_{2}>\ldots>X_{d}$
Integer $K\geq 0$ verifying the top-$K$ population rankings are correct with probability at least $1-\alpha$
$K\leftarrow 0$triangleright Number of verified ranks
while $K<d-1$ do
     $i\leftarrow K+1$triangleright Index of tested feature
     for $j\in[i+1:d]$ do
         Test $H_{0ij}:\{\mu_{i}<\mu_{j}\ |\ X_{i}>\max_{k>i}X_{k}\}.$triangleright Use original p-value $p_{ij}$ or SPRT statistic $T_{ij}^{n}$
     end for
     if all nulls $H_{0ij}$ reject then $\triangleright$ Equivalently, test $H_{0i}$ with $p_{i}^{*}\leftarrow\max_{j}p_{ij}$ or $T_{i*}^{n}\leftarrow\min_{j}T_{ij}^{n}$ (SPRT)
         $K\leftarrow K+1$
     else
         return $K$
     end if
end while
if $K=d-1$ then
     $K\leftarrow d$
end if
return $d$
Algorithm 4 Procedure 2 [Goldwasser et al., 2025]
Significance level $\alpha$, ordered data $X_{1}>X_{2}>\ldots>X_{d}$
Boolean whether top-$K$ population set is correct with probability at least $1-\alpha$
for $i\in[1:K]$ do
     for $j\in[K+1:d]$ do
         Test $H_{0ij}:\{\mu_{i}<\mu_{j}\ |\ X_{i}>\max_{k>K}X_{k}\}.$triangleright Use original p-value $p_{ij}$ or SPRT statistic $T_{ij}^{n}$
     end for
     if some null $H_{0ij}$ fails to reject then $\triangleright$ i.e. Test $H_{0i}$ with $p_{i}^{*}\leftarrow\max_{j}p_{ij}$ or $T_{i*}^{n}\leftarrow\min_{j}T_{ij}^{n}$ (SPRT)
         return False $\triangleright$ Cannot verify ith rank belongs in top-$K$ set
     end if
end for
return True
Appendix B Top-$K$ Set Algorithms
Algorithms 5 and 6 stabilize the top $K$ set. Like Algorithms 1 and 2, they employ the SPRT and resampling approaches, respectively.
Algorithm 5 Set Stability via SPRT
Desired set size $K>0$, error rate $\alpha\in[0,1]$, total sample budget $n_{max}$, number of samples between tests $n_{btwn}$, samples between tests $n_{btwn}$.
Estimated feature importances $\hat{\phi}_{1},\ldots,\hat{\phi}_{d}$ whose top-$K$ set is correct with probability $\geq 1-\alpha$.
$n\leftarrow 0$
while $n<n_{max}$ do
     Generate $n_{btwn}$ new samples
     $n\leftarrow n+n_{btwn}$
     $\hat{\phi}_{n}\leftarrow$ Feature importances fit on all $n$ samples
     $\hat{\Sigma}_{n}\leftarrow$ Variances of all feature importances
     Status $\leftarrow$ Procedure 2 (Alg. 4) on $\hat{\phi}_{n}$ $\hat{\Sigma}_{n}$, rejecting $H_{0ij}$ if $T_{ij}^{n}>\frac{1-\beta}{\alpha}$
     if Status==absent$=$ =“Verified" then
         return $\hat{\phi}$, “Verified"
     end if
end while
return $\hat{\phi}$, “Failed to verify"
Algorithm 6 Set Stability via Resampling
Desired rankings $K>0$, error rate $\alpha\in[0,1]$, per-feature sample budget $n_{max}$, intial per-feature samples $n_{init}$, buffer $c\geq 1$.
Estimated feature importances $\hat{\phi}_{1},\ldots,\hat{\phi}_{d}$ whose top-$K$ rankings are correct with probability $\geq 1-\alpha$.
ϕ^^italic-$\hat{\phi}\leftarrow$ Feature importances fit on $n_{init}$ samples
$\hat{\Sigma}\leftarrow$ Variances of all feature importances
while $n_{j}<n_{max}\;\forall j$ do
     Status $\leftarrow$ Procedure 2 (Alg. 4) on $\hat{\phi}_{n}$ $\hat{\Sigma}_{n}$, rejecting $H_{0ij}$ if $T_{ij}^{n}>\frac{1-\beta}{\alpha}$
     if Status==absent$=$ =“Verified" then
         return $\hat{\phi}$, “Verified"
     else
         $n_{i}^{\prime},n_{j}^{\prime}\leftarrow$ Estimated samples to reject, Eq. (10) or (11).
         $n_{i}^{\prime},n_{j}^{\prime}\leftarrow\min(\lceil cn_{i}^{\prime}\rceil,n_{max}),\min(\lceil cn_{j}^{\prime}\rceil,n_{max})$
         $\hat{\phi}_{i},\hat{\phi}_{j}\leftarrow$ Attributions fit on $n_{i}^{\prime},n_{j}^{\prime}$ samples
         $\hat{\Sigma}_{i},\hat{\Sigma}_{j}\leftarrow$ Variances of feature importances
     end if
end while
return $\hat{\phi}$, “Failed to verify"
Appendix C SPRT Proof
Here we prove Theorem 1, presenting the likelihood ratio necessary for SPRT. Returning to the terminology of Section 3.2, let $X_{j}\sim\mathcal{N}(\mu_{j},\sigma_{j}^{2})$ with independence and known $\sigma_{j}$; sort $X_{1}>X_{2}>\ldots X_{d}$.
Our objective is to verify feature $i$ as having the largest mean in a set of random variables, given it has the largest observed value $X_{i}$. Given a valid test to do so, Procedure 1 iterates from $i=1$ onwards until a failure to reject at $i=K+1$. At each step, $i$ is tested against all lower ranks. Procedure 2 conducts this test on $i=1$ through $K$ against features $j=K+1,\ldots,d$. Because rankings within the top-$K$ set do not matter, it only conditions on $\{X_{1}>\max_{j>K}X_{j}\}$.
Without loss of generality, let $i=1$. Define the null hypothesis
$H_{0}:$ $\mu_{1}\text{ {not} best}\ |\ X_{1}\text{ wins}$
[…]
Analogously define the alternate hypothesis
$H_{1}:$ $\mu_{1}\text{ {is} best}\ |\ X_{1}\text{ wins}$
[…]
A classical result states that a valid p-value for a union null hypothesis is the maximum p-value of its constituent nulls [Berger, 1982]. Therefore it suffices to construct valid tests for all $H_{0j}$; rejecting $H_{0}$ when all tests reject is a valid level-$\alpha$ procedure.
Ostensibly, SPRT tests $H_{0j}$ with a likelihood ratio $T_{1j}^{n}$, not a p-value. However, $T_{1j}^{n}$ may be interpreted probabilistically: Because SPRT is a valid test, the probability under the null that $T_{n}$ exceeds threshold $\gamma_{1}=\frac{1-\beta}{\alpha}$ is at most $\alpha$. Therefore the SPRT p-value is conceivably the probability of randomly obtaining a higher $T_{n}$ than the observed quantity, given that the null is true. Amongst all nulls $H_{0j}$, the p-value is highest for the lowest $T_{1j}^{n}$. So it suffices to compute $T_{1j}^{n}\;\forall j>1$, and reject $H_{0}$ if the smallest $T_{1j}^{n}\geq\gamma_{1}$. (Equivalently, $T_{1j}^{n}$ must exceed $\gamma_{1}$ for all $j$.)
Accepting the null is not a concern for reasonable $\alpha$ and $\beta$. For example, when $\alpha=0.05$ and $\beta=0.2$, the null is accepted when $T_{1j}^{n}$ is below roughly $0.21$. In practice, however, $T_{1j}^{n}$ should never even be below 1. Conditioning on the selection event that $X_{1}>X_{j}$, the data will always be likelier when the means mirror this discrepancy with $\Delta>0$.
Let $\Delta=\mu_{1}-\mu_{j}$. Under the null $\Delta\leq 0$, and under the alternate $\Delta>0$. Also let $A_{1}=\{X_{1}>\max_{k>1}X_{k}\}$, the event that $X_{1}$ wins. The $j$th likelihood ratio is then
$T_{1j}^{n}=\frac{\max_{\Delta>0}\mathbb{P}_{\mu_{1}-\mu_{j}=\Delta}(X_{1},\ldots,X_{d}\ |\ A_{1})}{\max_{\Delta\leq 0}\mathbb{P}_{\mu_{1}-\mu_{j}=\Delta}(X_{1},\ldots,X_{d}\ |\ A_{1})}.$
Following the same argument of Goldwasser et al. [2025], these probabilities may be simplified by conditioning on additional variables. The ensuing test will be valid at level-$\alpha$ for all possible realizations; as a result, the unconditional test would still be valid after marginalizing them out.
In particular, we condition on the values of non-tested variables $X_{-1j}$. We also condition on $U(X)=\frac{X_{1}}{\sigma_{1}^{2}}+\frac{X_{j}}{\sigma_{j}^{2}}$ taking its realized value, $u=\frac{x_{1}}{\sigma_{1}^{2}}+\frac{x_{j}}{\sigma_{j}^{2}}$. The purpose of this is to remove the influence of a nuisance parameter, which would otherwise prohibit inference on $\Delta$. We refer the reader to the manuscript for greater detail. The ratio that is tested instead is
$T_{1j}^{n}=\frac{\max_{\Delta>0}\mathbb{P}_{\mu_{1}-\mu_{j}=\Delta}(X_{1}\ |\ A_{1},X_{-1j},U)}{\max_{\Delta\leq 0}\mathbb{P}_{\mu_{1}-\mu_{j}=\Delta}(X_{1}\ |\ A_{1},X_{-1j},U)}.$ (12)
Ignoring constant factors, the conditional likelihood is proportional to the following:
$X_{1}\ |\ \{A_{1},X_{k\neq 1,j},U\}\propto\exp\left[-\left(\frac{1}{2\sigma_{1}^{2}}+\frac{\sigma_{j}^{2}}{2\sigma_{1}^{4}}\right)X_{1}^{2}+\left(\frac{\sigma_{j}^{2}u+\Delta}{\sigma_{1}^{2}}\right)X_{1}\right]\textbf{1}_{A_{1}}.$ (13)
Goldwasser et al. [2025] showed that this likelihood under the null is maximized at $\Delta=0$. The conditional distribution for $X_{1}$ that results is a truncated normal. Its parameters are defined in Section 3.2: Mean $\bar{\mu}_{1j}=\frac{\sigma_{j}^{2}x_{1}+\sigma_{1}^{2}x_{j}}{\sigma_{1}^{2}+\sigma_{j}^{2}}$, variance $\bar{\sigma}_{1j}^{2}=\frac{\sigma_{1}^{4}}{\sigma_{1}^{2}+\sigma_{j}^{2}}$, and truncation at $\bar{\eta}_{1j}=\max(\bar{\mu}_{1j},\max_{k\neq 1,j}x_{k})$. The denominator of (12) is thus
$\frac{\phi(\frac{x_{1}-\bar{\mu}_{1j}}{\bar{\sigma}_{1j}})\frac{1}{\bar{\sigma}_{1j}}}{1-\Phi(\frac{\bar{\eta}_{1j}-\bar{\mu}_{1j}}{\bar{\sigma}_{1j}})}.$ (14)
For arbitrary $\Delta$, the conditional likelihood is also a truncated normal. Its proportional density (13) may be rearranged to complete the square.
$X_{1}\ |\ \{A_{1},X_{k\neq 1,j},U\}$ $\propto\exp\left[\left(-\frac{\sigma_{1}^{2}+\sigma_{j}^{2}}{2\sigma_{1}^{4}}\right)X_{1}^{2}+\left(\frac{\sigma_{j}^{2}u+\Delta}{\sigma_{1}^{2}}\right)X_{1}\right]\textbf{1}_{A_{1}}$
$\propto\exp\left[-\frac{\sigma_{1}^{2}+\sigma_{j}^{2}}{2\sigma_{1}^{4}}\left(X_{1}^{2}-\frac{2\sigma_{1}^{2}(\sigma_{j}^{2}u+\Delta)}{\sigma_{1}^{2}+\sigma_{j}^{2}}X_{1}\right)\right]\textbf{1}_{A_{1}}$
$\propto\exp\left[-\frac{\sigma_{1}^{2}+\sigma_{j}^{2}}{2\sigma_{1}^{4}}\left(X_{1}-\frac{\sigma_{1}^{2}\sigma_{j}^{2}u+\sigma_{1}^{2}\Delta}{\sigma_{1}^{2}+\sigma_{j}^{2}}\right)^{2}\right]\textbf{1}_{A_{1}},$
This is a truncated normal distribution again with variance $\bar{\sigma}_{1j}^{2}$ and truncated at $\bar{\eta}_{1j}$. Its mean decomposes to the following:
$\tilde{\mu}_{1j}^{\Delta}=\frac{\sigma_{1}^{2}\sigma_{j}^{2}u+\sigma_{1}^{2}\Delta}{\sigma_{1}^{2}+\sigma_{j}^{2}}=\frac{\sigma_{j}^{2}x_{1}+\sigma_{1}^{2}x_{j}+\sigma_{1}^{2}\Delta}{\sigma_{1}^{2}+\sigma_{j}^{2}}.$
The numerator of Equation (12) is
$\max_{\Delta>0}\frac{\phi(\frac{x_{i}-\tilde{\mu}_{1j}^{\Delta}}{\bar{\sigma}_{1j}})\frac{1}{\bar{\sigma}_{1j}}}{1-\Phi(\frac{\bar{\eta}_{1j}-\tilde{\mu}_{1j}^{\Delta}}{\bar{\sigma}_{1j}})}.$ (15)
This expression cannot be manipulated into a closed-form solution for the optimal $\Delta$, due to the integral in the denominator. It could be optimized numerically, but doing so would be relatively slow. Furthermore, the SPRT algorithm must compute $T_{ij}^{n}$ for many $i,j$, and $n$, so this is impractical. Instead, a practical solution optimizes only the numerator of (15). The resulting choice of $\Delta$ may be slightly sub-optimal, producing a test that is barely more conservative.
$\Delta^{*}$ $=\operatorname*{arg\,max}_{\Delta>0}\phi(\frac{x_{i}-\tilde{\mu}_{1j}^{\Delta}}{\bar{\sigma}_{1j}})$
$=\operatorname*{arg\,max}_{\Delta>0}\exp\left[-\frac{(x_{1}-\tilde{\mu}_{1j}^{\Delta})^{2}}{2\bar{\sigma}_{1j}^{2}}\right]$
$=\operatorname*{arg\,min}_{\Delta>0}(x_{1}-\tilde{\mu}_{1j}^{\Delta})^{2}$
$=\operatorname*{arg\,min}_{\Delta>0}\left(\frac{\sigma_{1}^{2}x_{1}+\sigma_{j}^{2}x_{1}}{\sigma_{1}^{2}+\sigma_{j}^{2}}-\frac{\sigma_{j}^{2}x_{1}+\sigma_{1}^{2}x_{j}+\sigma_{1}^{2}\Delta}{\sigma_{1}^{2}+\sigma_{j}^{2}}\right)^{2}$
$=\operatorname*{arg\,min}_{\Delta>0}\left(\frac{\sigma_{1}^{2}}{\sigma_{1}^{2}+\sigma_{j}^{2}}(x_{1}+x_{j}-\Delta)\right)^{2}$
$=x_{1}-x_{j}.$
This is an intuitive choice: The optimal difference in means $\Delta^{*}$ is the observed difference, $x_{1}-x_{j}$. Furthermore, the conditional mean of $X_{1}$ under the alternate is the observed value $x_{1}$:
$\tilde{\mu}_{1j}^{\Delta^{*}}=\frac{\sigma_{j}^{2}x_{1}+\sigma_{1}^{2}x_{j}+\sigma_{1}^{2}x_{1}-\sigma_{1}^{2}x_{j}}{\sigma_{1}^{2}+\sigma_{j}^{2}}=\frac{\sigma_{j}^{2}x_{1}+\sigma_{1}^{2}x_{1}}{\sigma_{1}^{2}+\sigma_{j}^{2}}=x_{1}.$
Therefore the alternate likelihood (15) is
$\frac{\phi(0)\frac{1}{\bar{\sigma}_{1j}}}{1-\Phi(\frac{\bar{\eta}_{1j}-x_{1}}{\bar{\sigma}_{1j}})}.$
Composing this with (14) yields
$T_{1j}^{n}=\left[\frac{\phi(0)}{1-\Phi(\frac{\bar{\eta}_{1j}-x_{1}}{\bar{\sigma}_{1j}})}\right]\left[\frac{\phi(\frac{x_{i}-\bar{\mu}_{1j}}{\bar{\sigma}_{1j}})}{1-\Phi(\frac{\bar{\eta}_{1j}-\bar{\mu}_{1j}}{\bar{\sigma}_{1j}})}\right]^{-1}.$
Finally, the test statistic for $H_{0}$ is $\min_{j>1}T_{1j}^{n}$.
Appendix D Correlated Testing
The methodology we have presented assumes estimated ranks are independent. This assumption is necessary in order to apply the hypothesis tests in Equation (5) and Theorem 1. As described in Section 4, however, feature importances may be weakly correlated with one another. Violating this assumption is a limitation of our methodology. Fortunately, this limitation does not seem to affect practice, as the empirical FWERs of KernelSHAP are all controlled at level $\alpha$.
Nevertheless, it turns out that alternative tests can be constructed which take correlation into account. Unfortunately, they require making assumptions that are at least as concerning as indepedence itself.
To test each hypothesis $H_{0ij}$ (7), it is necessary to construct a density that isolates the parameter of interest, $\mu_{i}-\mu_{j}$. The joint density of all observations $X_{1:d}$ has a host of “nuisance parameters" — other unknown quantities that prohibit inference on $\mu_{i}-\mu_{j}$. To remove their inference, a requisite step conditions on the non-tested variables $X_{k}=x_{k}$, where $k\neq i,j$. When all variables are independent, the nuisance parameters $\mu_{k}$ disappear in the density of $X_{i,j}|X_{k\neq i,j}$. The remaining density can be rewritten in terms of $\mu_{i}-\mu_{j}$ and a final nuisance parameter, $\mu_{j}$.
When variables are correlated, the conditional distribution $X_{i,j}|X_{k\neq i,j}$ does not eliminate the nuisance parameters $\mu_{k\neq i,j}$. This is because the conditional means $\mathbb{E}[X_{i,j}\ |\ X_{k\neq i,j}]$ depend on $\mu_{k}$. To see this, let $X\sim\mathcal{N}(\mu,\Sigma)$, with indices $a=\{i,j\}$ and $b=[d]\backslash\{i,j\}$. Express the unconditional mean and variance as
$\mu=\begin{bmatrix}\mu_{a}\\ \mu_{b}\end{bmatrix}\text{\quad and \quad}\Sigma=\begin{bmatrix}\Sigma_{aa}&\Sigma_{ab}\\ \Sigma_{ba}&\Sigma_{bb}\end{bmatrix}.$
Classical results state the conditional distribution is normal with means and variance
$\mu_{a|b}$ $=\mu_{a}+\Sigma_{ab}\Sigma_{bb}^{-1}(x_{b}-\mu_{b}),$
$\Sigma_{a|b}$ $=\Sigma_{aa}-\Sigma_{ab}\Sigma_{bb}^{-1}\Sigma_{ba}$
Thus, the density of $X_{i,j}|X_{k\neq i,j}$ still contains all means $\mu_{k\neq i,j}$, vectorized in $\mu_{b}$. These nuisance parameters prohibit inference on $\mu_{i}-\mu_{j}$.
A strong assumption enables us to circumvent this restriction. The conditional means $\mu_{i}$ have their own ranking: Conditioned on $X_{k}$, we may compare $\mu_{i|b}$ and $\mu_{j|b}$. For all $i$ and $j$, we may assume the unconditional ordering of $\mu_{i}$ and $\mu_{j}$ is equal to the conditional ordering of $\mu_{i|b}$ and $\mu_{j|b}$. When this holds, a valid test of $H_{0ij}$ performs inference on the null that $\mu_{i|b}-\mu_{j|b}$.
It is possible to construct a test on these conditional means that takes their correlation $\rho_{ij}$ into account, using the conditional variance $\Sigma_{a|b}$. The argument to do so is identical to that of Goldwasser et al. [2025]:
  1. 1.
    Express the density in terms of the difference in means $\mu_{i|b}-\mu_{j|b}$ and a single nuisance parameter.
  2. 2.
    Conditioning on the ancillary statistic removes the nuisance parameter, and expresses the distribution in terms of $X_{i}$.
  3. 3.
    Apply the null hypothesis, under which $\mu_{i|b}-\mu_{j|b}=0$.
  4. 4.
    Completing the square reveals the mean and variance of a truncated normal distribution.
  5. 5.
    Identify the truncation event for $X_{i}$ winning in terms of $X_{k\neq i,j}$ and the ancillary statistic.
This approach may be followed for top-$K$ SPRT algorithms. The optimal densities under the null and alternative can be calculated, producing a likelihood ratio akin to Theorem (1).
We do not include these results and their proofs because the assumption of mean order preservation is very strong, and may be easily violated. For example, consider the simple three-variable case case with $\mu_{1}>\mu_{2}>\mu_{3}$. $X_{1}$ is independent of the other variables, but positive correlation exists between $X_{2}$ and $X_{3}$. Then whenever $X_{3}$ is observed to be relatively high, the conditional mean of $X_{2}$ is raised, perhaps higher than that of $X_{1}$.
In practice, it is impossible to know whether observed values are unusually high or low relative to their mean. Therefore this assumption is impossible to verify. Future work could explore conditions under which it holds. It is possible that this correlated test of conditional means is preferable to assuming independence, but further work is necessary to demonstrate its validity.
Appendix E Additional SHAP Experiments
E.1 StableSHAP Efficiency
We investigated how the adaptive allocation strategy of StableSHAP improved stability relative to Shapley Sampling. To do so, we compared their performance given equal computational budgets. The algorithms explained 30 test set predictions of a neural network trained on the Adult dataset. On each input, we ran StableSHAP (Alg. 2) to convergence 50 times, tracking the average number of total samples $n=\sum_{j}n_{j}$ (Equation (2)). Then, we ran Shapley Sampling 50 times, evenly distributing the same total number of samples across the $d$ features for each input. We repeated this procedure for $K=2$ and $K=5$, as well as $\alpha=0.1$ and $\alpha=0.2$.
Figure 5: Stability of StableSHAP and Shapley Sampling, Equal Computation.
Figure 5 displays empirical error rates on top-$K$ Shapley rankings. For all input samples, ranks $K$, and tolerances $\alpha$, StableSHAP successfully achieved FWER control, as anticipated. In contrast, rankings from Shapley Sampling were considerably less stable. In all settings, their error rates were higher than StableSHAP’s, often exceeding $\alpha$. This indicates that StableSHAP is more efficient than typical Shapley Sampling methods, which allocate samples equally across features.
E.2 Variances
Figure 6: Variances of StableSHAP values displayed in Figure 3
Figure 6 shows the variances of the SHAP estimates in Figure 3. They follow the same order by magnitude, with dashed line at $K=5$. While lower SHAP values generally have smaller variance, the 8th and 11th ranks have high variance. This accounts for the iterations in Figure 3 for which Occupation and Hours of Week require more than 100 samples.
Appendix F LIME Experiments
F.1 Methodology
LIME, introduced in Section 2.1.2, fits an interpretable model on data randomly generated around a point of interest. When linear models are used, their coefficients naturally are feature importance scores. For Ordinary Least Squares in particular, our retrospective and top-$K$ procedures may be applied directly.
For K-Lasso, in contrast, the $K$ features are selected iteratively along the Least-Angle Regression (LARS) path. Because data is regenerated at each step, our methods are not applicable in their current form. Rather, to ensure the FWER is controlled at level-$\alpha$, a multiple testing correction must be used. We apply the Bonferroni correction, performing each test at level $\alpha/K$. More sophisticated corrections like Holm’s method cannot be employed because $p$-values are observed sequentially [Holm, 1979]. Note this assumes LARS does not deselct any features, in which case more than $K$ tests would be performed.
At each step, LARS iteratively chooses the predictor that is most correlated with the current residuals. These scaled correlations $\hat{c}_{j}$ are asymptotically normal by the central limit theorem. In the retrospective case, verify whether the selected feature truly has highest correlation by performing Test (6) at level $\alpha/K$. To guarantee top-$K$ rank stability, run either the SPRT or resampling algorithm in Section 5 at level $\alpha/K$ until the top rank is stable.
We did not implement our algorithms on LIME from scratch. Rather, we repurposed an existing method, S-LIME, to adhere generally within our framework [Zhou et al., 2021]. S-LIME generates enough samples such that each highest-ranking feature beats its runner-up with probability exceeding $1-\alpha$, in a manner similar to Algorithm 2.
To verify the winner, Test (6) takes the highest p-value $p_{1j}$. Comparing to the runner-up is equivalent to a level-$\alpha/2$ Z-test. Unless all variances are equal, ranks below the runner-up may have higher p-values [Hung and Fithian, 2019]. To run S-LIME in this context, it is necessary to assume the runner-up always has the highest p-value.
Making the requisite assumptions, running S-LIME at level $\alpha/2K$ for each test is therefore a valid level-$\alpha$ ranking procedure for LIME with K-Lasso. The factors of 2 and $K$ account for selection and multiple testing, respectively.
F.2 Results
We ran the adjusted S-LIME procedure from Appendix F.1 on 5 datasets at $\alpha=0.1$ and $0.2$. As before, we compare 30 input samples, computing the error rate with 50 runs across each. Table 3 displays the experimental results.
For all inputs, the error rate is always controlled at level $\alpha$. In fact, this procedure is very conservative, with a maximal error rate of 4% for only one dataset. This may be attributed to its use of the Bonferroni correction, lowering the significance threshold by a factor of 5. Future experiments could run our top-$K$ algorithms on LIME with OLS, requiring no such correction.
Table 3: Maximum error rate (%) for top-$5$ ranking procedure on LIME with K-Lasso.
Dataset $K=2$ $K=2$ $K=5$ $K=5$
$\alpha=0.1$ $\alpha=0.2$ $\alpha=0.1$ $\alpha=0.2$
Adult 0% 2% NA NA
Bank 2% 0% 0% 0%
BRCA 2% 0% NA NA
Credit 6% 8% 0% 0%
WBC 2% 2% 0% 0%
Appendix G Experimental Details
For the models, we trained two-layer feedforward neural networks in Pytorch for 20 epochs. The hidden layer had 50 neurons. To discourage overfitting to the more common class, we batched the two classes in sampling each batch from the training data. We trained networks for 20 epochs in batches of 32 samples, using the Adam optimizer with a learning rate of 0.001.
Datasets for experimental results were pulled from the UCI Machine Learning repository. We random split the data 75%/25% into training and test sets using sklearn.
Computation was conducted using a slurm partition. Each entry in the tables was run with a separate job, though not with internal parallelization. In constructing Tables 2 and 3, we only considered input samples for which our algorithms successfully rejected all $K$ tests on at least 90% of the iterations.
G.1 SHAP
Our experiments ranked the absolute values of SHAP estimates. The sum of all SHAP values for input $x$ is $f(x)-E[f(X)]$, which may be positive or negative. In the latter case, the most important features are the ones with the largest negative values. Further, an individual sample may have features with both large positive and negative SHAP values. An alternative approach would be flip the sign of the SHAP values if their sum is negative, then rank these preprocessed values.
Because absolute values have non-negative support, their distribution is no longer Gaussian, but rather a folded normal. To avoid this, we used the same normal distribution, but flipped the sign if the SHAP estimate was negative. In theory, the true mean could then be the opposite sign as its observation. In essence, this does not consider the distribution of the absolute SHAP values. Instead, it merely establishes a ranking based on their absolute value. This subtle distinction may have virtually no practical difference, as the mass of the highest-ranking features may be almost entirely positive or negative.
Shapley Sampling (2) takes the average of $n$ values of $v(S_{j}^{i}\cup\{j\})-v(S_{j}^{i})$. To select subsets $S_{j}^{i}\subseteq[d]\backslash\{j\}$, each iteration randomly permutes the $d$ features, then takes the features that precede $j$. This is formalized in Algorithm 7.
Algorithm 7 Shapley Sampling
Input $x$, dataset $X$, Shapley value function $v(S)$, player $j\in[d]$, number of samples $n>0$
$\hat{\phi}_{j}(v)$, an unbiased estimate of $\phi_{j}(v)$
$\hat{\phi}_{j}\leftarrow 0$
for $i=1$ to $n$ do
     $\pi_{i}\leftarrow$ random permutation of $[d]$
     $S_{j}^{i}\leftarrow$ elements of $\pi_{i}$ before $j$
     $\hat{\phi}_{j}\leftarrow\hat{\phi}_{j}+v(S_{j}^{i}\cup\{j\})-v(S_{j}^{i})$
end for
$\hat{\phi}_{j}(v)\leftarrow\frac{1}{n}\hat{\phi}_{j}(v)$
To generate Figure 1, we used results from the experiment generating Figure 5. This compared StableSHAP and Shapley Sampling on the Adult Census Income dataset. We chose an input for which Shapley Sampling was less stable. On the same input, we ran SPRT-SHAP, LIME, and the adjusted S-LIME procedure described in Appendix F.1.
Our SHAP estimates computed $v(S)=\hat{\mathbb{E}}[f(X)|X_{S}=x_{s}]$ with 10 samples per subset $S$. Features were sampled from their marginal distributions.
We implemented KernelSHAP as described in Covert and Lee [2020]. The same paper introduces a method for estimating its variance. However, Goldwasser and Hooker [2023] found that a considerably more stable approach was to compute a bootstrap estimate. This method takes the covariance matrix of Shapley estimates fit on bootstrapped versions of the data. In our experiments, we used this approach with 250 bootstrapped samples.
Our retrospective experiments (Table 1) used $20482d+$ 2048 samples of subsets $S$. This is the default for KernelSHAP in the shap package. Note this is more than enough to guarantee convergence by CLT. Concerns of coarse approximation would only be legitimate under perhaps $n=30$ samples. For each of 30 input samples, the empirical FWER was computed over 50 runs. The table displays the maximum FWER over these 30 inputs.
Similarly, our top-$K$ experiments computed empirical FWERs over 50 runs, taking the maximum across 30 inputs. NA values in Table 2 correspond to settings of method, dataset, $K$, and $\alpha$ that proved incapable of converging with reasonable frequency. Specifically, we ceased computation after our methods failed to converge on fewer than 10% of inputs after at least 10 attempts.
StableSHAP used 100 initial samples per feature, and a maximum of 10,000. $n^{\prime}_{i}$ and $n^{\prime}_{j}$ were scaled according to their relative variances, following Equation (11). SPRT-SHAP iterations were capped at 50,000 samples. The SPRT test was conducted every 1,000 iterations. For the Type II error rate, we set $\beta=0.2$. This indicates the null will be accepted at most 20% of the time. (In practice, the null is never accepted.)
StableSHAP and SPRT-SHAP do not necessarily use the same inputs in Table 2. To construct Figure 4, we explicitly selected 30 inputs for which both top-$2$ ranking algorithms converged. For BRCA, there were not enough test set samples for which SPRT-SHAP stabilized the top-$2$ ranks under reasonable $\alpha$; therefore we used training samples on this dataset.
G.2 LIME
Our LIME procedure can be implemented using Zhengze Zhou’s slime repository almost entirely off-the-shelf. In our experiments, we added only 3 lines of code, which flag when the maximum number of samples have been used but not all hypothesis tests reject.
To control FWER at level $\alpha$, the “alpha" parameter passed to slime() should be $\frac{\alpha}{2K}$. We used 1,000 initial samples, and a maximum of 100,000. Another parameter, “tol," denotes the tolerance level of the hypothesis tests. Setting $tol=0$ corresponds to the algorithm in the paper, and is of course a viable option.
The creators of the package found that having a small positive tolerance could yield comparable results with considerably greater efficiency. In our experiments, we set $tol=10^{-4}$, which ran more quickly while controlling the FWER at level $\alpha$. While higher values of tol made S-LIME run more quickly, it also resulted in rare instances in which the algorithm allegedly converged without actually controlling the FWER. The arbitrary nature of this choice is a legitimate limitation of the S-LIME method. That said, our experimental errors are always below $\alpha$, and more a conservative approach could set $tol=0$.